# Sustainability Risks under Lotka-Volterra Dynamics

### v1

@@ 5 A generalised extension to an $N$-dimensional EVS model
The previous section provides a thorough analysis of the baseline 3D EVS system. The state variables take an aggregated form, and the interactive signs ($a_{ij}$) are set in a highly stylised manner in the baseline model. To take one further step, each state variable consists of many contributing factors, and these contributing factors may be inter-connected across the three dimensions or intra-connected within a particular dimension. We begin with the setup of an $N$-dimension Lotka-Volterra model that consists of $N=n_{1}+n_{2}+n_{3}$ variables, in which there are $n_{1}$ economic factors, $n_{2}$ environmental factors, and $n_{3}$ social factors, expressed in Eq. 36 - 39.
@@
The full analysis of the $N$-dimensional EVS model is beyond the scope of this paper, which will be further investigated in future research. There are, however, many existing studies in the field of applied mathematics on high-dimensional Lotka-Volterra systems, which may be relevant to understanding the full EVS model in our context. May (1972) demonstrates that for large linear systems with random coupling, the fixed points must jointly satisfy a fundamental inequality to ensure stability. May’s inequality remains a cornerstone of theoretical ecology, inspiring decades of research on complexity management in networks. Gouzé (1993) examines a generalised $N$-dimensional Lotka-Volterra (LV) model, which fundamentally transforms the analysis of global dynamics in the LV systems by reducing the problem to a matrix-algebraic framework. This approach rigorously establishes the universality of regular behaviour characterised by trajectories converging either to equilibria or diverging to infinity – across systems satisfying decomposition criteria. However, establishing the necessary and sufficient conditions for the existence of such decompositions remains an unresolved issue. Akjouj et al. (2024) propose an analytical framework based on a stochastic matrix for large-scale Lotka-Volterra models, which extends May (1972) and provides a rigorous mathematical foundation for understanding the stability of complex ecosystems. Future research may integrate network science with a machine learning approach to overcome analytical difficulties and computational bottlenecks in high-dimensional systems.
Another potentially interesting issue associated with the full model is the possible existence of chaos in a higher-dimensional model when $N>3$. Numerous studies investigate the chaotic dynamics in Lotka-Volterra systems. Vano et al. (2006) found that chaos can exist in a Lotka-Volterra model when the system consists of at least four dimensions. Christie et al. (2001) rigorously demonstrates, via Melnikov’s method, that periodic perturbations can induce Smale horseshoe chaos in Lotka-Volterra systems. This study establishes explicit parameter threshold conditions for chaotic behaviour and extends this framework to biomolecular self-organisation models, and establishes a novel paradigm for analysing complex dynamics in ecology and biophysics. The full mathematical properties of the generalised EVS system deserve further investigation in future research.
@@ 6 Conclusion
To this end, this paper proposes the EVS model: a generalised model of sustainability that characterises the interdependence between the economic ($E$), environmental ($V$), and social ($S$) factors in alignment with the sustainability definition of WCED (1987). The EVS model is inspired by earlier works of Meadows et al. (1972), Meadows et al. (1992), Herrington (2021) and Todorov and Marinova (2011). Our model has several advantages that can address the earlier critics to the World3 model. First, the model has a better analytical tractability thanks to its simplistic mathematical structure based on a Lotka-Volterra setting. Despite its simplicity, the system can fairly characterise the competitive and cooperative relationship amongst the state variables of our interests. Second, we adopt a Schumpetarian definition of economic development to better characterise the innovation-induced sustainability risks, compared with a relatively static assumption of technological development in the World3 model. Third, the model has a high degree of flexibility, as it can be decomposed into multiple subdynamics to address specific sustainability risks within a particular sector, as well as to establish connections between the subsystem and other components, thereby understanding sustainability risks from a holistic perspective. The model can also shed light on the boundary between the private and public sectors in maintaining sustainable development. While the private sector is driven by Schumpeterian-type disruptive innovation and technological dissemination, sustainability risks can arise, which require the public sector to constantly monitor its sustainability risks and ensure that the control parameters remain confined to the sustainability regime by enforcing better environmental regulation, encouraging innovation in the circular economy, and fostering social development that ultimately serves the economy and the environment.
@@
The current version of our EVS model is still in its infancy, which can be extended and applied in numerous directions, both empirically and theoretically. It is also important to note, from a practical perspective, that the EVS framework can be applied to the sustainability risk assessment of emerging technologies in a systematic manner. While ESG can still serve as a means for promoting corporate sustainability, EVS can be applied to assess the risks associated with existing and disruptive technologies, particularly regarding their role in aggravating or mitigating climate risks, as well as other environmental hazards to prevent apocalyptic consequences beyond the point of control. These extensions and applications will be further investigated in future research.
Appendix A Mathematical Proof of Proposition 1
Condition 1 (Social sustainability)
When the EV subsystem approaches its equilibrium point $(E^{*},V^{*})$,
$\frac{dS}{dt}\bigg{|}_{(E,V)\to(E^{\star},V^{\star})}>0,$ (40)
Condition 2 (Environmental sustainability)
When the ES subsystem approaches its equilibrium point $(E^{*},S^{*})$,
@@ Appendix A Mathematical Proof of Proposition 1
Condition 3 (Economic sustainability)
When the VS subsystem approaches its nodal equilibrium $(V^{*},S^{*})$,
@@
To satisfy Condition 1, we set $S=0$. The system is reduced to an EV subdynamics:
which yields
$\left\{\begin{aligned} V^{*}&=-\frac{r_{E}}{a_{12}},\\ E^{*}&=-\frac{r_{V}}{a_{21}}.\end{aligned}\right$. (44)
Furthermore,
$\frac{dS}{dt}=S(r_{S}+a_{31}E+a_{32}V)>0.$ (45)
Hence,
$r_{S}>\frac{a_{32}r_{E}}{a_{12}}+\frac{a_{31}r_{V}}{a_{21}}=\frac{a_{32}a_{21}r_{E}+a_{31}a_{12}r_{V}}{a_{21}a_{12}}.$ (46)
According to our Assumptions,
$a_{21}<0,a_{12}>0,a_{32}>0,r_{S}<0\ and\ r_{E}<0,$
which yields
$a_{21}a_{12}r_{S}-a_{32}a_{21}r_{E}<a_{31}a_{12}r_{V}.$
Then,
[…]
Similarly, based on Condition 2, we set $V=0$. The system is reduced to an ES subdynamics:
@@
$\left\{\begin{aligned} S^{*}&=-\frac{r_{E}}{a_{13}},\\ E^{*}&=-\frac{r_{V}}{a_{31}}.\end{aligned}\right$. (49)
Condition 2 indicates that
@@
[…]
According to our assumptions,
$r_{V}>0,r_{E}<0,a_{31}<0,a_{13}>0,a_{23}>0.$
Therefore, $a_{13}r_{V}>0>\ a_{23}r_{E}$ always holds, implying
$a_{31}<\frac{a_{12}a_{13}r_{S}}{a_{13}r_{V}-a_{23}r_{E}}.\quad$ (52)
To satisfy Condition 3, we set $E=0$. The VS subdynamics is set as follows:
@@
$\left\{\begin{aligned} S^{*}&=-\frac{r_{V}}{a_{23}},\\ V^{*}&=-\frac{r_{S}}{a_{32}},\end{aligned}\right$. (54)
$\frac{\mathrm{d}E}{\mathrm{d}t}=E(r_{E}+a_{12}V^{*}+a_{13}S^{*})>0.$ (55)
According to our assumptions,
$a_{12}>0,\quad a_{13}>0,$ (56)
which yields
[…]
Q.E.D.
@@ Appendix B Numerical Simulation of the 3D Baseline EVS Model
The authors are listed in alphabetical order, and both authors have made equal contributions to this paper. This project is initiated by Tianhao Zhi, following his long-term research interests in sustainable development and nonlinear economic dynamics. Tianhao has contributed to the initial idea of applying a Lotka-Volterra model to sustainability risks, conducted a substantial part of the literature review, outlined the modelling structure, and wrote the first draft of this paper and the MATLAB code for numerical simulations. He presented the preliminary version of this paper at the International Day for Biodiversity 2024 cum Guangdong-Hong Kong-Macao Greater Bay Area Nature Education and Ecological Economic Forum organised by Dr Siu-Tai Tsim, Associate Head of the Department of Life Sciences, Beijing Normal-Hong Kong Baptist University. At a later stage, Tianhao Zhi invited Yiren Wang to join the project after a period of regular discussion with each other on this topic. Yiren has made a substantial contribution to surveying the technical literature regarding recent developments in the Lotka-Volterra model, formulated the mathematical definition of sustainability, provided the mathematical derivations and proofs of the 3D model, and suggested the potential application with a stochastic matrix in the $N$-dimensional scenario. The remaining errors are on our own.

### v2

@@ 5 A generalised extension to an $N$-dimensional EVS model
The previous section provides a thorough analysis of the baseline 3D EVS system. The state variables take an aggregated form, and the interactive signs ($a_{ij}$) are set in a highly stylised manner in the baseline model. To take one further step, each state variable consists of many contributing factors, and these contributing factors may be inter-connected across the three dimensions or intra-connected within a particular dimension - for example, the creative destruction process can be manifested as a competitive LV relationship within the $E$ subdimension between the new technology and the old technology. At the same time, the development of quantum computing and AI can be characterised by a cooperative LV relationship. We set up an $N$-dimensional Lotka-Volterra model that consists of $N=n_{1}+n_{2}+n_{3}$ variables, in which there are $n_{1}$ economic factors, $n_{2}$ environmental factors, and $n_{3}$ social factors, expressed in Eq. 33 - 36.
@@
The full analysis of the $N$-dimensional EVS model is beyond the scope of this paper, which will be further investigated in future research. There are, however, many existing studies in the field of applied mathematics on high-dimensional Lotka-Volterra systems, which may be relevant to understanding the full EVS model in our context. May (1972) demonstrates that for large linear systems with random coupling, the fixed points must jointly satisfy a fundamental inequality to ensure stability. May’s inequality remains a cornerstone of theoretical ecology, inspiring decades of research on complexity management in networks. Gouzé (1993) examines a generalised $N$-dimensional Lotka-Volterra (LV) model, which fundamentally transforms the analysis of global dynamics in the LV systems by reducing the problem to a matrix-algebraic framework. This approach rigorously establishes the universality of regular behaviour characterised by trajectories converging either to equilibria or diverging to infinity – across systems satisfying decomposition criteria. However, establishing the necessary and sufficient conditions for the existence of such decompositions remains an unresolved issue. Akjouj et al. (2024) propose an analytical framework based on a stochastic matrix for large-scale Lotka-Volterra models, which extends May (1972) and provides a rigorous mathematical foundation for understanding the stability of complex ecosystems. Future research may further look at the empirical estimation of the interactive matrix in the EVS system to understand the full scope of competitive or cooperative relationships among the $N$ variables. This can potentially be done by integrating a machine learning approach to overcome analytical difficulties and computational bottlenecks in high-dimensional systems.
Another potentially interesting issue associated with the full model is the possible existence of chaos in a higher-dimensional model when $N>3$. Numerous studies investigate the chaotic dynamics in Lotka-Volterra systems. Vano et al. (2006) found that chaos can exist in a Lotka-Volterra model when the system consists of at least four dimensions. Christie et al. (2001) rigorously demonstrates, via Melnikov’s method, that periodic perturbations can induce Smale horseshoe chaos in Lotka-Volterra systems. This study establishes explicit parameter threshold conditions for chaotic behaviour and extends this framework to biomolecular self-organisation models, and establishes a novel paradigm for analysing complex dynamics in ecology and biophysics.
Last but not least, the recent developments in dynamical mean field theory (DMFT) would also be relevant to the full analysis of the EVS system. DMFT is developed and applied to analyse dynamic systems consisting of a large number of interacting agents by reducing it to a one-dimensional master equation in terms of mean-field interaction. Its application in social sciences is popularised by Weidlich and Haag (1983). More recently, Roy et al. (2019) adopts DMFT to an $N$-dimensional Lotka-Volterra model, which is relevant in our context of an $N$-dimensional EVS system. The full mathematical properties of the generalised EVS system deserve further investigation in future research.
@@ 6 Conclusion
While global warming is an insidious process driven by accumulated fuel consumption over the past century, disruptive technologies emerging today may pose sustainability risks with much greater social and ecological impacts over a shorter time horizon. It is therefore vital to systematically monitor and manage sustainability risks at an early stage of the technological dissemination to prevent unforeseen social and ecological disasters. To this end, this paper proposes the EVS model: a generalised model of sustainability that characterises the interdependence between the economic ($E$), environmental ($V$), and social ($S$) factors in alignment with the sustainability definition of WCED (1987). The EVS model connects all the state variables in the three sub-dimensions through their competitive and cooperative relationships, and thus can be applied to understand the sustainability risks of disruptive technology in a systematic manner. The EVS model is inspired by earlier works of Meadows et al. (1972), Meadows et al. (1992), Herrington (2021) and Todorov and Marinova (2011). Our model has several advantages that can address the earlier criticisms of the World3 model. First, the model has a better analytical tractability thanks to its simplistic mathematical structure based on a Lotka-Volterra setting. Despite its simplicity, the system can fairly characterise the competitive and cooperative relationship amongst the state variables of our interests. Second, we adopt a Schumpetarian definition of economic development to better characterise the innovation-induced sustainability risks, rather than the relatively static assumption of technological development in the World3 model. Third, the model has a high degree of flexibility, as it can be decomposed into multiple subdynamics to address specific sustainability risks within a particular sector, as well as to establish connections between the subsystem and other components, thereby understanding sustainability risks from a holistic perspective. The model can also shed light on the boundary between the private and public sectors in maintaining sustainable development. While the private sector is driven by Schumpeterian-type disruptive innovation and technological dissemination, sustainability risks can arise, which require the public sector to constantly monitor and manage its sustainability risks and ensure that the control parameters remain confined to the sustainability regime by enforcing better environmental regulation, encouraging innovation in the circular economy, and fostering social development that ultimately serves the economy and the environment.
@@
The current version of our EVS model is still in its infancy, which can be extended and applied in numerous directions, both empirically and theoretically. While ESG can still serve as a means to promote corporate sustainability, EVS can be applied to evaluate sustainability risks associated with the existing and emerging technologies, particularly regarding their roles in exacerbating or mitigating climate risks, as well as other environmental and social risks, to prevent catastrophic consequences beyond the point of control. These extensions and applications will be further investigated in future research.
Appendix A The fixed point resilience for the ES and VS subdynamics
Condition 1 (Environmental resilience of the ES subdynamics)
When the ES subsystem approaches its fixed point $(E^{*},S^{*})$,
@@ Appendix A The fixed point resilience for the ES and VS subdynamics
Condition 2 (Economic resilience of the VS subdynamics)
When the VS subsystem approaches its fixed point $(V^{*},S^{*})$,
@@
To determine the environmental sustainability condition, we set $V=0$. The system is reduced to an ES subdynamics:
@@
$\left\{\begin{aligned} S^{*}&=-\frac{r_{E}}{a_{13}},\\ E^{*}&=-\frac{r_{S}}{a_{31}}.\end{aligned}\right$. (40)
This fixed point may be unattainable since the sign of $a_{31}$ is undetermined. In case if $a_{31}>0$, both $S^{\star}$ and $E^{\star}$ take positive signs. The environmental sustainability condition requires that:
@@
$a_{31}(a_{13}r_{V}-a_{23}r_{E})>a_{21}a_{13}r_{S}.$ (42)
Since, $r_{V}>0$, $r_{E}<0$, $a_{31}>0$, $a_{13}>0$, and $a_{23}>0$, it implies that $a_{13}r_{V}-a_{23}r_{E}>0$. We have:
$a_{31}>\frac{a_{21}a_{13}r_{V}}{a_{13}r_{V}-a_{23}r_{E}}.$ (43)
Similarly, to determine the economic sustainability condition, we set $E=0$. The VS subdynamics is set as follows:
@@
$\left\{\begin{aligned} S^{*}&=-\frac{r_{V}}{a_{23}},\\ V^{*}&=-\frac{r_{S}}{a_{32}}.\end{aligned}\right$. (45)
This fixed point is unattainable since $r_{V}>0$ and $a_{23}>0$, which implies that $S^{\star}<0$.
@@ Appendix B Numerical Simulation of the 3D Baseline EVS Model
The authors are listed in alphabetical order, and both authors have made equal contributions to this paper. This project is initiated by Tianhao Zhi, following his long-term research interests in sustainable development and nonlinear economic dynamics. Tianhao has contributed to the initial idea of applying a Lotka-Volterra model to sustainability risks, conducted a substantial part of the literature review, outlined the modelling structure, and wrote the first draft of this paper and the MATLAB code for numerical simulations. He presented the preliminary version of this paper at the International Day for Biodiversity 2024 cum Guangdong-Hong Kong-Macao Greater Bay Area Nature Education and Ecological Economic Forum organised by Dr Siu-Tai Tsim, Associate Head of the Department of Life Sciences, Beijing Normal-Hong Kong Baptist University. At a later stage, Tianhao Zhi invited Yiren Wang to join the project after a period of regular discussion with each other on this topic. Yiren has made a substantial contribution by surveying the technical literature regarding recent developments in the Lotka-Volterra model, formulated the mathematical definition of sustainability, and provided the mathematical derivations and proofs of the 3D model. The remaining errors are on our own.
@@ Acknowledgement
  • 2.
    The authors would like to thank Howard Huxter for introducing the two works of the Club of Rome to Tianhao Zhi a few years ago: Limits to Growth and Mankind at the Turning Point. It serves as an inspiration for the Tianhao’s continuing research on sustainable development.
  • 3.
    The authors would like to thank participants at the International Day for Biodiversity 2024 cum Guangdong-Hong Kong-Macao Greater Bay Area Nature Education and Ecological Economic Forum in Zhuhai, China on 18 May 2024, the 2025 WEHIA (Winter) Workshop of Economics with Heterogeneous Interacting Agents at XJ-Liverpool University in Suzhou, China (13 - 16 February 2025), and the 2025 Risk Sciences Annual Conference at the Nanyang Technological University in Singapore on 10 September 2025 for valuable feedback and suggestions.
  • 4.
    The usual disclaimer applies.